OpenStudy (anonymous):

Simplify (4x − 5)^2.

OpenStudy (anonymous):

(a+b)^2=a^2 + b^2 + 2ab

OpenStudy (anonymous):

does that mean "multiply"? start with \[(4x+5)(4x+5)\] and do four multiplications

OpenStudy (anonymous):

(a-b)^2 = a^2 + b^2 -2ab

OpenStudy (anonymous):

oops rather \[(4x-5)(4x-5)\]

OpenStudy (anonymous):

thankss(:

OpenStudy (anonymous):

do four multiplications, combine like terms, and then tell your teacher (gently) that there is nothing "simpler" about your result than there is in \((4x-5)^2\)

OpenStudy (anonymous):

is it 16x+25?

OpenStudy (anonymous):

oh no! careful here

OpenStudy (anonymous):

four multiplications i will write them out

OpenStudy (anonymous):

\[(4x-5)(4x-5)=4x\times 4x-5\times 4x-5\times 4x+5\times 5\]

OpenStudy (anonymous):

(4x-5)^2 = 16x^2 -40x +25.

OpenStudy (anonymous):

16x-40x+25?

OpenStudy (anonymous):

then \(4x\times 4x=16x^2\) \(-5\times 4x-5\times 4x=-20x-20x=-40x\) \((-5)\times (-5)=25\) final answer is \[16x^2-40x+25\]

OpenStudy (anonymous):

now factorize it or solve by formula \[x=-b \pm \sqrt{b^{2}-4ac}/2a\]

OpenStudy (anonymous):

no marmar it's 16x^2 - 40x +25. not 16x

OpenStudy (anonymous):

now use quadratic formula or apply factorization method.it's up to u

OpenStudy (anonymous):

U got it marmar? just use formula (a-b)^2 = a^2 + b^2 -2ab. then find x by factorization or quadratic formula.Simple!

OpenStudy (anonymous):

yea, thanks guys!